# acmart.bib

### v1

@@ Abstract.
We propose a novel training-free scoring mechanism to evaluate QML circuit performance on imbalanced classification and regression tasks. Our mechanism demonstrates significant correlation between scoring metrics and test performance on imbalanced classification tasks. Additionally, we develop methods to quantify continuous similarity relationships between quantum states, enabling performance prediction for regression tasks. This represents the first comprehensive approach to searching and evaluating QCS circuits specifically for regression applications.
@@
Validation on representative ADME tasks—one imbalanced classification and one regression—demonstrates moderate positive correlation between our scoring metrics and circuit performance, significantly outperforming baseline scoring methods that show negligible correlation.
@@
Quantum computing, Quantum Circuit Search, Drug property prediction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computer systems organization Quantum computingccs: Applied computing Life and medical sciencesccs: Computing methodologies Machine learning
Figure 1. Overview of QCS-ADME
1. Introduction
@@
Quantum Machine Learning (QML) has gained significant attention for its potential applications on Noisy Intermediate-Scale Quantum (NISQ) hardware (Preskill_2018), demonstrating promising results across various domains (Huang_2021; Huang_2021_E; silver2023mosaiqquantumgenerativeadversarial; Huang_2022; Liu_2021). QML architectures differ fundamentally from classical machine learning models in their use of data embedding gates, where gate selection and placement significantly impact performance, and in their gradient computation methods, which employ techniques like parameter-shift rules (Schuld_2019) rather than traditional backpropagation. These unique characteristics make manual circuit design particularly challenging, often resulting in suboptimal performance.
To address this challenge, Quantum Circuit Search (QCS) (Zhang_2022; Wang_2022; Du_2022) has emerged as a methodology for discovering high-performance, noise-robust circuits, drawing inspiration from classical Neural Architecture Search (NAS) (li2019randomsearchreproducibilityneural). State-of-the-art QCS approaches (Wang_2022; 10.1145/3620665.3640354) generate circuit candidates by employing diverse algorithmic designs and evaluating performance metrics, taking into account hardware specifications and constraints.
The biomedical field has begun exploring QML applications for tasks traditionally addressed by classical machine learning, particularly in ADME (absorption, distribution, metabolism, and excretion) property prediction—crucial characteristics in drug evaluation. However, ADME tasks present unique challenges for existing QCS frameworks, as they encompass both classification and regression problems. Current QCS methods, even the Elivagar (10.1145/3620665.3640354), a specialized QCS method designed for QML tasks, designed primarily for classification tasks on balanced datasets like MNIST, are not directly applicable to ADME’s frequently imbalanced datasets and regression requirements.
@@
To address these limitations, we propose two key methodological enhancements: a weighted matrix approach for unbalanced classification tasks that amplifies minority class representation, and a modified representational capacity score for continuous similarity comparison in regression tasks. We validate these improvements on representative ADME tasks—one unbalanced classification problem and one regression problem—demonstrating improved correlation between our enhanced scoring methods and test performance compared to traditional approaches.
In summary, our contributions are:
1. We propose the first application of the Quantum Circuit Synthesis (QCS) framework to generate quantum machine learning (QML) circuits for biomedical ADME tasks (Absorption, Distribution, Metabolism, Excretion).
2. We introduce a novel scoring mechanism for the QCS framework to evaluate circuit performance on imbalanced classification and regression tasks, addressing key challenges in biomedical data analysis.
3. We demonstrate that QML circuits achieve competitive advantages over classical models in specific scenarios (e.g., lower mean squared error in regression tasks) and identify actionable pathways to further improve their robustness for ADME applications.
2. Method
@@
To address the requirements of ADME Property Prediction, which encompasses both regression tasks and classification tasks with imbalanced datasets, we enhanced the QCS framework mainly in two key aspects: circuit performance evaluation on imbalanced datasets and regression task handling. Specifically, we chose the Elivagar (10.1145/3620665.3640354), a state-of-the-art system originally designed for searching QML circuits in balanced binary classification tasks, as our baseline model to improve. This section first outlines Elivagar’s original circuit scoring methodology, followed by our proposed improved structure to accommodate imbalanced classification and regression tasks.
2.1. Overview of Elivagar Scoring
@@
Elivagar’s circuit evaluation process consists of two main stages: circuit generation and scoring. Initially, the system generates candidate circuits based on device topology and noise-related parameters. Each candidate circuit is then evaluated using two predictive metrics: Clifford Noise Resilience (CNR) and Representational Capacity (RepCap). CNR predicts circuit fidelity based on structural characteristics, while RepCap quantifies the discriminative power of the circuit by measuring both intra-class similarities and inter-class separation in the final quantum states. These scores are combined to produce a final metric, where higher values indicate greater potential for superior performance.
@@
2.2. Handling Class Imbalance
@@
(1) $\text{RepCap}(C)=1-\frac{\|R_{c}-R_{ref}\|^{2}_{2}}{2\cdot n_{c}\cdot d_{c}^{2}}$
Here, $d_{c}$ denotes the number of samples and $n_{c}$ represents the number of classes. $R_{ref}$ is a $d\times d$ reference matrix representing ideal circuit predictions, where $R_{ref}(i,j)=1$ if samples $i$ and $j$ belong to the same class ($y_{i}=y_{j}$) and 0 otherwise. The matrix $R_{c}$ is also of dimension $d\times d$ and quantifies the similarity between quantum states of data points $i$ and $j$. While this formulation performs effectively on balanced datasets, it exhibits limitations when handling imbalanced class distributions. In balanced scenarios, the reference matrix contains equal representation for each class. However, when certain classes have significantly more samples than others, the RepCap metric may assign high scores to circuits that excel at distinguishing majority classes while performing poorly on minority classes. To address this limitation, we incorporate classical imbalanced classification techniques by introducing a weight matrix to the reference matrix, thereby amplifying the importance of minority classes. This modification ensures that circuits must develop robust representations for all classes, regardless of their distribution in the training data, to achieve high scores. The enhanced RepCap formulation is:
(2) $\text{RepCap}(C)=1-\frac{\|R_{c}-R_{w}\otimes R_{ref}\|^{2}_{2}}{2\cdot n_{c}\cdot d_{c}^{2}},$
where $R_{w}$ represents the weight matrix derived from the class distribution in the training data, and $\otimes$ denotes the element-wise multiplication operation.
2.3. Regression Task Adaptation
@@
(3) $R_{ref}=\exp\left(-\frac{\|y_{i}-y_{j}\|_{2}^{2}}{2\sigma^{2}}\right),$
where $\|y_{i}-y_{j}\|_{2}^{2}$ represents the squared Euclidean distance between target values, and $\sigma$ serves as a bandwidth parameter controlling the decay rate of similarity with respect to distance.
The second fundamental modification addresses the quantum state similarity computation. While retaining the basic quantum state similarity calculation framework, we incorporate continuous weighting based on target values. This adaptation is essential because conventional quantum state similarity measurements only capture state relationships in Hilbert space, without considering the proximity of corresponding target values. By introducing target value-dependent Gaussian weighting factors, we ensure the quantum circuit learns to map inputs with similar target values to correspondingly similar quantum states.
@@
3. Experiments
3.1. Dataset Selection and Preprocessing
From the TDC (velez-arce2024signals) ADME group, we selected two representative tasks: Cell Effective Permeability (Caco2) (Wang2016ADME) for regression and Parallel Artificial Membrane Permeability Assay (PAMPA) (doi:10.1177/24725552211017520) for binary classification. The Caco2 dataset requires predicting cell effective permeability values from drug SMILES strings. The PAMPA dataset involves classifying compounds as either high permeability (1) or low-to-moderate permeability (0) based on their SMILES strings. Notably, the PAMPA dataset exhibits significant class imbalance, with a ratio of approximately 6:1 between high and low permeability samples.
To prepare these datasets for our QML circuits, we implemented several preprocessing steps. First, we transformed the SMILES strings into Molecular ACCess Systems keys fingerprints (MACCS) (doi:10.1021/ci010132r). MACCS fingerprints are 166-bit strings where each bit indicates the presence (1) or absence (0) of specific molecular structural features. We utilized the first 128 bits to accommodate our quantum system’s requirements. For the classification task, we remapped all 0 labels to -1. For the regression task, we normalized the target values in both training and testing sets to the range [-1, 1].
3.2. Training and Evaluation Configuration

### v2

# QCS-ADME: Quantum Circuit Search for Drug Property Prediction with Imbalanced Data and Regression Adaptation

@@ Abstract
We propose a novel training-free scoring mechanism to evaluate QML circuit performance on imbalanced classification and regression tasks. Our mechanism demonstrates significant correlation between scoring metrics and test performance on imbalanced classification tasks. Additionally, we develop methods to quantify continuous similarity relationships between quantum states, enabling performance prediction for regression tasks. This represents a novel training-free approach to searching and evaluating QCS circuits specifically for regression applications.
@@
Validation on representative ADME tasks—eight imbalanced classification and four regression—demonstrates moderate correlation between our scoring metrics and circuit performance, significantly outperforming baseline scoring methods that show negligible correlation.
keywords:
@@
Quantum machine learning, Quantum circuit search, Drug Property Prediction
Figure 1: Overview of QCS-ADME workflow. The pipeline proceeds from A) converting SMILES string to classical bitstring, to B) mapping classical bitstring to qubit space via quantum embeeding, to C) generate candidate circuits and compute their fidelity and RepCap metrics, to D) Select the top circuits and train them.
1 Introduction
@@
Quantum Machine Learning (QML) has gained significant attention for its potential applications on Noisy Intermediate-Scale Quantum (NISQ) hardware [Preskill_2018], demonstrating promising results across various domains [Huang_2021, Huang_2021_E, silver2023mosaiqquantumgenerativeadversarial, Huang_2022, Liu_2021]. QML architectures differ fundamentally from classical machine learning models in their use of data embedding gates, where gate selection and placement significantly impact performance, and in their gradient computation methods, which employ techniques like parameter-shift rules [Schuld_2019] rather than traditional backpropagation. These unique characteristics make manual circuit design particularly challenging, often resulting in suboptimal performance.
Quantum Circuit Search (QCS) [Zhang_2022, Wang_2022, Du_2022], which is inspired from classical Neural Architecture Search (NAS) [li2019randomsearchreproducibilityneural], has emerged as a methodology for discovering high-performance, noise-robust circuits to tackle the above challenges. State-of-the-art QCS approaches [Wang_2022, 10.1145/3620665.3640354] generate circuit candidates by employing diverse algorithmic designs and evaluating performance metrics, taking into account hardware specifications and constraints. However, these methods have largely been benchmarked on simple, balanced datasets such as MNIST and Moons, limiting their applicability to complex, real-world problems. This limitation is particularly acute in the biomedical field, where predicting ADME properties is crucial for drug evaluation. ADME tasks present unique challenges that existing QCS frameworks: they frequently involve highly imbalanced classification datasets and complex regression targets.
While advancements like Quantum-SMOTE [Mohanty2025] offer data-level solutions for imbalance, they rely on computationally expensive quantum swap tests. Within a QCS framework which requires rapid evaluation of thousands of candidates, such approaches introduce computation overhead. Consequently, there is currently no QCS scoring mechanism capable of effectively handling the dual challenges of regression topology and class imbalance found in pharmacological data.
@@
To bridge this gap, we propose QCS-ADME, a novel framework that adapts QCS for biomedical property prediction. We introduce two key methodological enhancements: a weighted matrix approach for unbalanced classification tasks that amplifies minority class representation, and a modified representational capacity score for continuous similarity comparison in regression tasks. We validate these improvements on representative ADME tasks—eight unbalanced classification problems and four regression problem—demonstrating improved correlation between our enhanced scoring methods and test performance compared to traditional approaches.
In summary, our work not only adopts the fundamental search strategy from previous QCS methods, but also mitigates the limitations that make previous QCS methods ineffective for biomedical ADME tasks. Detailed novelty and contributions are listed as follows:
1. We develop a novel training-free scoring mechanism capable of handling regression tasks. Unlike prior methods that maximize class separability using discrete matching, our Gaussian-weighted RepCap generalizes the metric from discrete to continuous space. This encourages the quantum circuit to learn a geometry in Hilbert space that is isomorphic to the continuous topology of regression targets.
2. We introduce a density-aware weighting matrix $R_{w}$ to address class imbalance. This formulation modifies the search landscape to penalize circuits that collapse onto majority classes, correcting a critical failure point where standard QCS metrics yield high scores despite poor minority class performance.
3. We provide the application of QCS to ADME property prediction. Furthermore, we conduct a comprehensive sim-to-real analysis, quantifying the performance gap between noiseless simulation and real hardware execution.
2 Background
2.1 Quantum Circuit Search (QCS)
Quantum Circuit Search (QCS) is a systematic approach that uses machine learning or optimization-driven techniques to automate the design of quantum circuits. Inspired by its classical counterpart, Neural Architecture Search (NAS), any QCS framework can be broken down into three core components: a search space that defines all possible circuits, a search strategy to navigate this space, and a performance evaluation method to assess the quality of candidate circuits. The search strategy is the engine of a QCS framework. Three dominant paradigms have emerged in this domain: Reinforcement Learning approach [patel2024curriculumreinforcementlearningquantum, Zhu2023] frames circuit design as a sequential decision-making process where a software agent learns a policy for placing gates to maximize a cumulative reward signal; Evolutionary Algorithms [e25010093, Zhang2023, e27070733] approach evolve circuit populations through genetic operators like crossover and mutation; Differentiable Search [HE2024106508, chen2024differentiablequantumarchitecturesearch] relaxes discrete gate choices into continuous parameters amenable to gradient-based optimization.
The single greatest obstacle in QCS is the prohibitive cost of performance evaluation. Fully training each candidate circuit to assess its quality is computationally intractable. One plausible paradigm is training-free QCS [10.1145/3620665.3640354]. This approach bypasses costly training entirely by using efficiently computable proxies to predict a circuit’s performance. Instead of measuring the final outcome, these proxies analyze a circuit’s intrinsic properties that are precursors to good performance, such as its expressibility, its topological complexity, or the features of its cost function landscape.
2.2 ADME Properties
For a small-molecule drug to be effective and safe, it must successfully navigate a complex journey through the body. This journey is characterized by a set of pharmacokinetic properties known as ADME, which stands for Absorption, Distribution, Metabolism, and Excretion. These properties sequentially dictate how a drug is first absorbed into the bloodstream from its site of administration, then distributed from the blood to various tissues to reach its site of action. Subsequently, the body metabolizes the drug, chemically breaking it down, before it is finally excreted along with its metabolites.
A drug candidate must possess a balanced and favorable ADME profile to exert its therapeutic effect without causing toxicity. In fact, a poor ADME profile is one of the most prominent reasons for the failure of drug candidates during expensive and lengthy clinical trials. Therefore, the ability to accurately predict a compound’s ADME properties from its chemical structure early in the drug discovery process is crucial. Early and reliable ADME profiling allows researchers to prioritize candidates with a higher likelihood of success, saving significant time and resources. [velez-arce2024signals]
3 Method
@@
To address the requirements of ADME Property Prediction, which encompasses both regression tasks and classification tasks with imbalanced datasets, we enhanced the QCS framework mainly in two key aspects: circuit performance evaluation on imbalanced datasets and regression task handling. The overall workflow of our proposed QCS-ADME framework is illustrated in Figure 1. The pipeline consists of four distinct phases:
  • A. Data Preprocessing: The pipeline begins by converting chemical structures into SMILES strings, which are then encoded into 166-bit MACCS keys. These are further processed to form a fixed-length 128-bit binary input vector.
  • B. Quantum Feature Mapping: The classical 128-bit input vector is mapped into the quantum Hilbert space using an Angle Embedding Layer, resulting in a 7-qubit feature vector.
  • C. Quantum Circuit Search: We evaluate a pool of candidate circuits using a composite scoring metric. This involves calculating the Clifford Noise Resilience (CNR) for hardware efficiency and the Representational Capacity (RepCap). The RepCap calculation incorporates our novel weighted matrix $R_{w}$ for imbalanced classification and Gaussian similarity for regression tasks. The final ranking score is computed as Score = $\text{CNR}^{\alpha_{\text{CNR}}}\times\text{RepCap}$.
  • D. Execution and Evaluation: The top-$k$ circuits with the highest scores are selected for execution. The parameters $\theta$ are iteratively updated via a classical optimizer to minimize the prediction loss.
The structure of this section is as follow: We first outlines Elivagar’s original circuit scoring methodology, then we introduce our proposed improved structure to accommodate imbalanced classification and regression tasks.
3.1 Circuit Generation Scoring
@@
The circuit evaluation process consists of two main stages: circuit generation and scoring. Initially, the system generates candidate circuits based on device topology and noise-related parameters. We follow the basic framework from Elivagar [10.1145/3620665.3640354], each candidate circuit is evaluated using two predictive metrics: Clifford Noise Resilience (CNR) and Representational Capacity (RepCap). CNR is a training-free proxy for circuit fidelity that predicts noise robustness primarily from a circuit’s structural characteristics (e.g. depth, entangling pattern, two-qubit gate exposure) by evaluating efficiently simulable Clifford replicas of the candidate circuit, while RepCap quantifies the discriminative power of the circuit by measuring both intra-class similarities and inter-class separation in the final quantum states. These scores are combined to produce a final metric, where higher values indicate greater potential for superior performance.
@@
3.2 Handling Class Imbalance
@@
$\text{RepCap}(C)=1-\frac{\|R_{c}-R_{ref}\|^{2}_{2}}{2\cdot n_{c}\cdot d_{c}^{2}}$ (1)
Here, $d_{c}$ denotes the number of samples and $n_{c}$ represents the number of classes. $R_{ref}$ is a $d\times d$ reference matrix representing ideal circuit predictions, where $R_{ref}(i,j)=1$ if samples $i$ and $j$ belong to the same class ($y_{i}=y_{j}$) and 0 otherwise. The matrix $R_{c}$ is also of dimension $d\times d$ and quantifies the similarity between quantum states of data points $i$ and $j$. Formally, we define the elements of $R_{c}$ as the fidelity between the variational states in Hilbert space, which constitutes the empirical Quantum Kernel $K_{Q}$:
$R_{c}(i,j)\equiv K_{Q}(x_{i},x_{j})=|\langle\psi(x_{i})|\psi(x_{j})\rangle|^{2}$ (2)
By minimizing the Frobenius distance between this kernel matrix and the ideal reference $R_{ref}$, RepCap maximize the alignment between kernel and the target. This ensures the selected circuit naturally maps data into a geometry that reflects the class structure, reducing the burden on the classical optimizer and improving generalization potential.
While this formulation performs effectively on balanced datasets, it exhibits limitations when handling imbalanced class distributions. In balanced scenarios, the reference matrix contains equal representation for each class. However, when certain classes have significantly more samples than others, the RepCap metric may assign high scores to circuits that excel at distinguishing majority classes while performing poorly on minority classes. To address this limitation, we incorporate classical imbalanced classification techniques by introducing a weight matrix to the reference matrix, thereby amplifying the importance of minority classes. This modification ensures that circuits must develop robust representations for all classes, regardless of their distribution in the training data, to achieve high scores. The enhanced RepCap formulation is:
$\text{RepCap}(C)=1-\frac{\|R_{c}-R_{w}\otimes R_{ref}\|^{2}_{2}}{2\cdot n_{c}\cdot d_{c}^{2}},$ (3)
We define the weight matrix $R_{w}$ as $R_{w}(i,j)=w_{i}\cdot w_{j}$, where $w_{i}$ is the inverse class frequency associated with sample $i$. Specifically, $w_{i}=\frac{N}{N_{c}}$, where $N$ is total samples and $N_{c}$ is the count of samples in class $c$. This ensures that pairs of minority-class samples contribute significantly more to the RepCap score, forcing the circuit to distinguish them.
3.3 Regression Task Adaptation
@@
$R_{ref}=\exp\left(-\frac{\|y_{i}-y_{j}\|_{2}^{2}}{2\sigma^{2}}\right),$ (4)
where $\|y_{i}-y_{j}\|_{2}^{2}$ represents the squared Euclidean distance between target values, and $\sigma$ serves as a bandwidth parameter controlling the decay rate of similarity with respect to distance. This formulation effectively treats $R_{ref}$ as a Radial Basis Function (RBF) kernel on the labels, encoding a smoothness prior common in kernel regression: pairs with similar targets should remain close in the representation space. Consequently, aligning the quantum similarity matrix $R_{c}$ with the RBF label kernel $R_{ref}$ promotes a specific geometry where increasing Euclidean distance in the label space corresponds to increasing Fubini distance in the Hilbert space.
The second modification addresses the quantum state similarity computation. While retaining the basic quantum state similarity calculation framework, we incorporate continuous weighting based on target values. This adaptation is essential because conventional quantum state similarity measurements only capture state relationships in Hilbert space, without considering the proximity of corresponding target values. By introducing target value-dependent Gaussian weighting factors, we ensure the quantum circuit learns to map inputs with similar target values to correspondingly similar quantum states.
@@
4 Experiments
4.1 Dataset Selection and Preprocessing
From the TDC [velez-arce2024signals] ADME and Tox groups, we selected eight classification tasks: Human Intestinal Absorption (HIA_Hou) [hou2007adme], P-glycoprotein Inhibition (Pgp_Broccatelli) [doi:10.1021/jm101421d], Bioavailability (Bioavailability_Ma) [MA2008677], Blood-Brain Barrier (BBB_Martins) [martins2012bayesian], CYP2C9/CYP2D6 Substrate (CYP2C9_Substrate_CarbonMangels, CYP2D6_Substrate_CarbonMangels) [https://doi.org/10.1002/minf.201100069], Human ether-à-go-go related gene blockers (hERG) [doi:10.1021/acs.molpharmaceut.6b00471], Drug Induced Liver Injury (DILI) [doi:10.1021/acs.jcim.5b00238]. And four regression tasks: Volumn of Distribution at steady state (VDss_Lombardo) [doi:10.1021/acs.jcim.6b00044], Half Life of a drug (Half_Life_Obach) [obach2008trend], Drug clearance (Clearance_Hepatocyte_AZ, Clearance_Microsome_AZ) [DI2012441]. Note that the classification tasks here have an imbalance positive to negative ratio from 3:1 to 6:1.
To prepare these datasets for our QML circuits, we implemented several preprocessing steps. First, we transformed the SMILES strings into Molecular ACCess Systems keys fingerprints (MACCS) [doi:10.1021/ci010132r]. MACCS fingerprints are 166-bit strings where each bit indicates the presence (1) or absence (0) of specific molecular structural features. We utilized the first 128 bits to accommodate our quantum system’s requirements. For the classification task, we remapped all 0 labels to -1. For the regression task, we normalized the target values in both training and testing sets to the range [-1, 1].
4.2 Quantum Circuit Methodology
Figure 2: One candidate quantum circuit found for the BBB_Martins task. The circuit is transpiled for the IBM Rensselaer processor, utilizing 7 qubits with angle encoding and variational layers containing parameterized rotations and hardware-native ECR entangling gates.
Circuit Architecture and Encoding: We employ device-aware variational quantum circuits optimized for the 127-qubit IBM Rensselaer superconducting processor. For the ADME tasks, we utilize a subsystem of $n=7$ qubits. Given a molecular descriptor vector $x\in\mathbb{R}^{128}$, we employ an Angle Encoding strategy where each feature $x_{j}$ parameterizes a single-qubit rotation. Specifically, the data embedding layer consists of a sequence of $R_{Y}$ rotations:
$R_{Y}(x_{j})|0\rangle=\cos\left(\frac{x_{j}}{2}\right)|0\rangle+\sin\left(\frac{x_{j}}{2}\right)|1\rangle$ (5)
The full circuit integrates 128 such data-encoding gates interleaved with variational parameters, resulting in a dense encoding of the feature space. The ansatz contains 32 trainable parameters optimized during the learning process. Figure 2 illustrates a representative optimized circuit architecture found by our search algorithm. This diagram explicitly visualizes the native gate decomposition used during execution on the hardware.
Noise-Aware Circuit Construction: Candidate circuits for the search space are generated via a biased random sampling strategy that incorporates real-time device calibration data. The sampling probability for placing a gate on specific qubits is weighted by:
  • Qubit coherence times ($T_{1}$, $T_{2}$)
  • Two-qubit gate fidelities (CNOT/ECR error rates)
  • Measurement readout fidelities
  • Physical hardware connectivity constraints
This probability distribution biases the construction toward higher-fidelity qubits and links, effectively pruning the search space of hardware-inefficient architectures before training begins.
Measurement and Prediction: The final prediction is derived from the expectation value of the Pauli-$Z$ operator, measured on a single readout qubit $q_{out}$ selected for optimal readout fidelity:
$f(x,\theta)=\langle\psi(x,\theta)|Z_{q_{out}}|\psi(x,\theta)\rangle\in[-1,1]$ (6)
For binary classification tasks, the discrete label is assigned via $\operatorname{sign}(f(x,\theta))$.
4.3 Training and Evaluation Configuration